1 NUMBER THEORY

By Habel Mathewkutty .

Polyhedrons are Geometrical shapes enclosed by polygons. Numbers on them can be represented by Habel Math formula Akn = 2k(n-1) + 1 Habelmath sum = Hkm = (m/3)k(m - 1)(2m - 1) + 6 2+ 2(k +1)+ 2(4k + 1)+ 2(9k +1) + 2(16k +1)+.........+ 2(m-1)k + 1=H where H = (m/3)k(m-1)(2m - 1) + 6 Habel Math’s wonderful formula for sum to m terms of all Polyhedral numbers. Remember k = 1 for Tetrahedron, and k=29 for Soccerball because we know the soccerball numbers are A29n = 229(n - 1) + 1 They are 2, 60, 234, 524, ................... So H29m = (m/3)29(m - 1)(2m - 1) + 6 When m=4 it should be 2+60+234+524 = 820 By Prof. Habel Mathewkutty M. Sc.(Math/Agra), Ph. D. Speaker of SIAM conference NW08 in Rome 21-24 July 2008. Former Researcher of Indian Institutes of Technology and Instructor of Houston Community College System.